# When Text Embedding Meets Large Language Model: A Comprehensive Survey

### v2

@@ Abstract
Text embedding has become a foundational technology in natural language processing (NLP) during the deep learning era, driving advancements across a wide array of downstream tasks. While many natural language understanding challenges can now be modeled using generative paradigms and leverage the robust generative and comprehension capabilities of large language models (LLMs), numerous practical applications—such as semantic matching, clustering, and information retrieval—continue to rely on text embeddings for their efficiency and effectiveness. Therefore, how to combine the LLMs and the text embeddings has become one of the hotspots of academic attention in recent years. In this survey, we categorize the interplay between LLMs and text embeddings into three overarching themes: (1) LLM-augmented text embedding, enhancing traditional embedding methods with LLMs; (2) LLMs as text embedders, adapting their innate capabilities for high-quality embedding; and (3) Text embedding understanding with LLMs, leveraging LLMs to analyze and interpret embeddings. By organizing recent works based on interaction patterns rather than specific downstream applications, we offer a novel and systematic overview of contributions from various research and application domains in the era of LLMs. Furthermore, we highlight the unresolved challenges that persisted in the pre-LLM era with pre-trained language models (PLMs) and explore the emerging obstacles brought forth by LLMs. Building on this analysis, we outline prospective directions for the evolution of text embedding, addressing both theoretical and practical opportunities in the rapidly advancing landscape of NLP.
@@ 1 Introduction
Text embedding learning is a fundamental task in Natural Language Processing (NLP), which aims to extract features from text with the natural languages (including words, sentences, and documents). Formally, given the text $\mathcal{X}$, text embedding learning aims at training a model $f:\mathcal{X}\times\mathrm{\theta}\rightarrow\mathbb{R}^{d}$ on the dataset $D\subset\mathcal{X}$. The dense vector $h=f_{\theta}(x)$ output by $f$ is called the embedding of the text $x$, which is expected to contain valid information in the input text and perform well on the downstream tasks.
Advanced large language models (LLMs) have recently demonstrated exceptional generalization capabilities in downstream tasks where generative paradigms are applicable, such as information extraction, text classification, and machine translation. However, not all NLP tasks are suitable for modeling with generative paradigms; tasks such as dense retrieval, semantic text similarity, etc., still require text embedding for computing similarity. The powerful semantic understanding capability of the LLMs reveals the ability to use it to obtain high-quality embeddings. However, decoder-only LLMs are not different from traditional encoder-only pre-trained language models (PLMs), facing the anisotropy challenge in the native embedding space [1]. It is manifested by the high similarity of two token embeddings, which does not reflect their semantic similarity. Higher-level text, such as sentences and documents, suffer from similar problems since they often share embedding space with tokens [2]. Fortunately, LLMs have comprehension and generation capabilities far beyond previous PLMs, which opens up new opportunities for text-embedding learning. Specifically, LLMs change the existing landscape in two ways: (1) LLMs can act as data annotators or generators for large-scale, high-quality, and fine-grained text datasets, and (2) LLMs can replace current PLMs as the backbone for generating higher-quality text embeddings. Many works have been devoted to introducing LLMs to obtain high-quality text embedding; however, they focus on a single downstream task, and similar methods have been proposed repeatedly in different research fields. This leads to a lack of comprehensive and objective understanding of the role played by LLMs in text embedding research.
@@
Figure 1: The overview of the five relationships between text embedder and LLMs. The correspondence between symbols and meanings is as follows: $\mathcal{I}$: Instruction, $\mathcal{E}$: Example, $\mathcal{X}$: Text, $\mathcal{H}$: Text Embedding, $\mathcal{Y}$: Label.
@@ 2.1.1 Era of Statistics Machine Learning
Text embedding is first applied in information retrieval, text mining, and machine translation. In the early days, researchers primarily relied on manually designed features to represent text, which are used to measure the relation between the texts. These methods require domain experts to select and design features carefully, and the quality and quantity of features limit their effectiveness. The development of the method is accompanied by a transformation of the output vector’s form, which is from bag-of-words models (one-hot vectors) to TF-IDF [6] (sparse vectors), then to text embeddings (dense vectors). As machine learning techniques advanced, researchers began exploring methods to learn text embeddings using statistical approaches such as Latent Semantic Analysis (LSA) [7] for information retrieval and Latent Dirichlet Allocation (LDA) [8] for topic mining. Though these methods can automatically learn low-dimensional text embedding, there are still obvious shortcomings. For example, LSA is hard to capture complex semantic and syntactic structures, while LDA is difficult to apply to large-scale datasets.
@@ 2.1.4 Era of Pre-trained Language Models
where $s:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the distance function, while $x^{+}$ and $\{x^{-}_{j}\}_{j=1}^{N}$ is the positive example and negative examples for the text $x$, respectively. Improvements based on the contrastive learning have focused on (a) the form of the loss function [81, 82, 83, 84] and (b) better methods for constructing positive and negative samples [79, 85, 86, 80]; (3) Pre-training paradigm. Contrastive learning is useful on large-scale corpora but takes big performance hits on low data situations [74]. This observation shows that the pre-training tasks of PLMs, i.e., Masked Language Modeling (MLM) and Next Sentence Prediction (NSP), etc., do not allow the PLMs to adapt to semantic aggregation quickly and output high-quality text embeddings. Therefore, some work has been devoted to designing better pre-training tasks to improve the performance of contrastive learning in low-data scenarios. Early explorations [87, 88] used only additional incremental pre-training tasks, such as the Inverse Cloze Task (ICT), to improve contrast learning performance. Subsequent improvements involve the modification of model architectures. For example, Condenser [89] restores the masked token using the [CLS]’s hidden states from the last layer and other tokens’ hidden states from a shallow layer. Therefore, the newly generated information in the deep layers has to aggregate in the [CLS] to reconstruct the masked information. SEED-Encoder [90] and RetroMAE [91] feed the final hidden state of [CLS] into an auxiliary decoder to restore the masked information, encouraging the aggregation of information at [CLS].
@@ 2.3.1 Semantic Textual Similarity
Semantic Textual Similarity (STS) task [109] is the task to judge the similarity calculated by embeddings whether matching the ratings from humans. Strictly speaking, STS is not a downstream task with a well-defined application scenario, but rather an intrinsic evaluation task that was purposely presented [110]. The motivation behind it is that the distance between two texts’ embedding should reflect the semantics similarity between these two texts.
@@
The dataset for evaluating STS can be expressed as $D_{\rm eval}^{\rm STS}=\{(x_{i}^{a},x_{i}^{b},c_{i})\}_{i=1}^{n}$, where $(x_{i}^{a},x_{i}^{b})$ is a text pair and $c_{i}$ is the average similarity between the two texts. The popular evaluation datasets usually contains seven STS datasets for English setting, which contains STS12-16 [109, 111, 112, 113, 114], STS-B [110] and SICK-R [115] in SentEval Benchmark [116] and STS-17 [110] and STS-22 [117] for multi-lingual setting.
@@
For evaluation, all texts $\{x_{i}^{a}\}_{i=1}^{n}\cup\{x_{i}^{b}\}_{i=1}^{n}$ are encoded by the learned embedder. And a similarity function, e.g. dot product or cosine similarity, is introduce to measure the semantic similarity between the text pairs in the embedding space. Common evaluation metrics include Pearson correlation coefficient and Spearman correlation coefficient.
@@
Since different methods lead to large differences in the value range of similarity, and we are concerned with relative rankings rather than absolute values in practice [118]. Therefore, Spearman correlation coefficient is used by default.
@@ 2.3.2 Information Retrieval
Information retrieval (IR)  [^7] Information retrieval in this survey refers to ad-hoc text retrieval by default. aims to retrieve the most related texts relevant to the query from a large-size candidate sets. Modern IR systems involves multiple stages of recalling or re-ranking from different-scale candidate sets. Embedding-based retrieval methods are referred to as dense retrieval, which is often used as a method of recalling or secondary-filtering, placed after sparse retrieval (e.g., BM25 [119]) and before reranking (e.g., the cross-encoder scheme recommended by BERT  [68]).
@@
For efficiency of evaluation, all query $\{q_{i}\}_{i=1}^{n}$ and all candidate texts $\{d_{i}\}_{i=1}^{n}\cup\{d_{j}\}_{j=1}^{N}$ can be encoded as embeddings with the learned embedder in advance. Then a similarity function is introduced to measure the relevance for each query-candidate embedding pair. For each query, all candidate texts will sorted by the predicted relevance in descending and form a ordered candidate list. The common evaluation metrics include Recall Rate, Accuracy, Mean Reciprocal Rank (MRR), Mean Average Precision (MAP) and Normalized Discounted Cumulative Gain (NDCG).
@@
    where $\text{R}_{q_{i}}^{(1)}$ is the rank of the frist relevant text in the ordered candidate list.
@@
  • NDCG considers the rank of the relevant text, and suggest situating the more pertinent text at a more superior position. DCG for each $q_{i}$ should be calculated first, which can be written as
@@
    where $r_{i}$ is the graded relevance score for the $i$-th retrieved text in the candidae text list. Then $\text{NDCG}@k$ can be calculated by aggregating the normalized DCG values at a specific rank position:
@@ 2.3.3 Universal Embedding
The MTEB benchmark [122] contains 56 datasets in 7 different types of downstream tasks, including classification, clustering, pair classification, reranking, retrieval, and Semantic Textual Similarity (STS), to assess the degree of generalization of text embedding. Until October 2024, MTEB has the community variants in Chinese [123], French [124], Polish [125], Scandinavian [126], Russian [127], Arabic [128] and Persian [129], and a official multilingual version, i.e. MMTEB  [^8] https://github.com/embeddings-benchmark/mteb/blob/main/docs/mmteb, during developing.
@@ 2.4.2 Embedding Inversion
The main metrics used for evaluating embedding inversion include BLEU, ROUGE, Exact-Match, cosine similarity, and Token F1. Please refer to Equation 12 and 13 for Exact-Match and ROUGE-L, separately.
  • BLEU is used to measure n-gram similarity between reconstructed text and original input text. we can compute it in the following:
@@ 4 LLMs as Text Embedder
LLMs commonly adopt encoder-decoder architecture, e.g., T5 [94], or decoder-only architecture, e.g., GPT-3 [162]. After going through pre-training, instruction tuning, and alignment with human preferences [163], LLMs can perform a wide range of tasks. However, these steps do not support obtaining high-quality embeddings from LLMs. To address this issue, exploring methods to obtain text embeddings from LLMs directly is a promising direction. In Table 4, we list 37 LLM-based Embedders and their detailed information. Specifically, we focus on their backbone selection, architectural improvement, training methods, and evaluation protocols.
@@ 4.1 Backbone Selection
The LLM-based embedders appearing in Table 4 use numerous open-source LLMs as backbones. Among the encoder-decoder backbones, T5 (GTR) has a dominant position, which is not surprising, as T5 and its variants have been leading innovations in encoder-decoder PLMs. Among decoder-only backbones, the most popular is Mistral (17 / 37), followed by LLaMA (15 / 37) and Qwen (5 / 37). The popularity of the Mistral is strongly connected to its excellent performance: a recent study [164] find that Mistral’s embedding performance will not drop significantly compared to the other LLMs when converting causal attention to bi-directional attention. This characteristic may derive from its pre-training or instruction fine-tuning phase, but no information has been revealed.
@@
Interestingly, the only several attempts to use LLM with larger LLMs lead to even worse performance [165, 166].
@@ 4.2.1 Pooling Strategy
PromptEOL [166] OPT, LLaMA Last Causal $\times$ $\times$ Prompt TF ZS / ICL STS / Clf.
@@
Token Prepending [173] LLaMA, Qwen, Gemma Last Causal $\times$ $\times$ - TF ZS / FT STS / Clf.
BeLLM [174] LLaMA Last Bi $\times$ LoRA - SCL ZS / FT STS / Clf.
@@
PromptReps [180] Mistral, Phi-3, LLaMA Last Causal $\surd$ $\times$ Instruction+Prompt TF ZS IR
LMORT [181] GPT2, GPT-J Multi-Layer Causal $\times$ $\times$ - SCL FT IR
@@
RARe [183] LLaMA, Mistral Mean / Last Bi / Causal $\times$ LoRA Instruction+Examples SCL ICL IR
@@
Echo [185] Mistral Partial Mean Causal $\times$ LoRA Instruction+Prompt TF ZS Uni.
@@
LLM2Vec [164] LLaMA, Mistral Mean Bi $\times$ LoRA Instruction MNTP→UCL ZS Uni.
@@
BGE-ICL [196] Mistral, Gemma Special Last Bi / Causal $\times$ LoRA Instruction+Examples SCL ZS / ICL Uni.
@@
  • Special-Token Last Pooling Some works [189, 190, 176, 180, 179, 178, 177] introduce a special token, such as <EOS>, to obtain embedding in the last position, while the LLM is fine-tuned incrementally, learning to converge semantics at the position of the special token. These methods are marked as “special last” in the “pooling” column of Table 4.
  • Special-Sequence Last Pooling ChatRetriever [179] and DEBATER [184] add a special token sequence, such as $[\texttt{EMB}_{1}],\cdots,[\texttt{EMB}_{t}]$, at the end of the input text. The authors argue that these $t$ consecutive special tokens serve as a chain of thought that extends and guides the learning space for more effective embeddings. These methods are marked as “special seq last” in the “pooling” column of Table 4.
@@
Some works add complex modules with attention mechanisms before the conventional pooling strategy. NV-Embed [193] introduces an additional attention-based network before mean pooling to post-interact with the hidden states $\mathcal{H}$. The attention-based network consists of a cross-attention block and an MLP, where the key and value matrix used in the cross-attention are learnable and low-rank, and $\mathcal{H}$ is considered as the query. Mistral-SPLADE [182] uses the prompt same to Echo [185] to obtain text embeddings and follow the SPLADE recipe [198] to fine-tune the Mistral-based embedder.

### v3

@@ Abstract
Text embedding has become a foundational technology in natural language processing (NLP) during the deep learning era, driving advancements across a wide array of downstream tasks. While many natural language understanding challenges can now be modeled using generative paradigms and leverage the robust generative and comprehension capabilities of large language models (LLMs), numerous practical applications—such as semantic matching, clustering, and information retrieval—continue to rely on text embeddings for their efficiency and effectiveness. Therefore, integrating LLMs with text embeddings has become a major research focus in recent years. In this survey, we categorize the interplay between LLMs and text embeddings into three overarching themes: (1) LLM-augmented text embedding, enhancing traditional embedding methods with LLMs; (2) LLMs as text embedders, adapting their innate capabilities for high-quality embedding; and (3) Text embedding understanding with LLMs, leveraging LLMs to analyze and interpret embeddings. By organizing recent works based on interaction patterns rather than specific downstream applications, we offer a novel and systematic overview of contributions from various research and application domains in the era of LLMs. Furthermore, we highlight the unresolved challenges that persisted in the pre-LLM era with pre-trained language models (PLMs) and explore the emerging obstacles brought forth by LLMs. Building on this analysis, we outline prospective directions for the evolution of text embedding, addressing both theoretical and practical opportunities in the rapidly advancing landscape of NLP.
@@ 1 Introduction
Text embedding learning is a fundamental task in Natural Language Processing (NLP), aiming to extract features from text at various levels, including words, sentences, and documents. Formally, given the text $\mathcal{X}$, text embedding learning aims at training a model $f:\mathcal{X}\times\mathrm{\theta}\rightarrow\mathbb{R}^{d}$ on the dataset $D\subset\mathcal{X}$. The dense vector $h=f_{\theta}(x)$ output by $f$ is called the embedding of the text $x$, which is expected to contain valid information in the input text and perform well on the downstream tasks.
Advanced large language models (LLMs) have recently demonstrated exceptional generalization capabilities in downstream tasks where generative paradigms are applicable, such as information extraction, text classification, and machine translation. However, not all NLP tasks are suitable for modeling with generative paradigms; tasks such as dense retrieval, semantic text similarity, etc., still require text embedding for computing similarity. The powerful semantic understanding capability of the LLMs reveals the ability to use it to obtain high-quality embeddings. However, like traditional encoder-only pre-trained language models (PLMs), decoder-only LLMs face the anisotropy challenge in their native embedding space [1]. It is manifested by the high similarity of two token embeddings, which does not reflect their semantic similarity. Higher-level text, such as sentences and documents, suffer from similar problems since they often share embedding space with tokens [2]. Fortunately, LLMs have comprehension and generation capabilities far beyond previous PLMs, which opens up new opportunities for text-embedding learning. Specifically, LLMs change the existing landscape in two ways: (1) LLMs can act as data annotators or generators for large-scale, high-quality, and fine-grained text datasets, and (2) LLMs can replace current PLMs as the backbone for generating higher-quality text embeddings. Many works have been devoted to introducing LLMs to obtain high-quality text embedding; however, they focus on a single downstream task, and similar methods have been proposed repeatedly in different research fields. This leads to a lack of comprehensive and objective understanding of the role played by LLMs in text embedding research.
@@
Figure 1: An overview of the five relationships between text embedders and LLMs. The correspondence between symbols and meanings is as follows: $\mathcal{I}$: Instruction, $\mathcal{E}$: Example, $\mathcal{X}$: Text, $\mathcal{H}$: Text Embedding, $\mathcal{Y}$: Label.
@@ 2.1.1 Era of Statistics Machine Learning
Text embedding is first applied in information retrieval, text mining, and machine translation. In the early days, researchers primarily relied on manually designed features to represent text, which are used to measure the relation between the texts. These methods require domain experts to carefully select and design features, and their effectiveness is constrained by both the quality and quantity of these features. The development of the method is accompanied by a transformation of the output vector’s form, which is from bag-of-words models (one-hot vectors) to TF-IDF [6] (sparse vectors), then to text embeddings (dense vectors). As machine learning techniques advanced, researchers began exploring methods to learn text embeddings using statistical approaches such as Latent Semantic Analysis (LSA) [7] for information retrieval and Latent Dirichlet Allocation (LDA) [8] for topic mining. Although these methods can automatically learn low-dimensional text embeddings, they still have significant shortcomings. For example, LSA is hard to capture complex semantic and syntactic structures, while LDA is difficult to apply to large-scale datasets.
@@ 2.1.4 Era of Pre-trained Language Models
where $s:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the distance function, while $x^{+}$ and $\{x^{-}_{j}\}_{j=1}^{N}$ is the positive example and negative examples for the text $x$, respectively. Improvements based on the contrastive learning have focused on (a) the form of the loss function [81, 82, 83, 84] and (b) better methods for constructing positive and negative samples [79, 85, 86, 80]. For example, ANCE [87] explores a two-stage training method. First, it utilizes non-embedding model methods, such as BM25, to extract a batch of negative samples for model training. Subsequently, it employs the trained embedding model to re-mine higher-quality negative samples for further model training. This method has become a standard practice for training embedding models. Conan-Embedding [88] employs a dynamic hard negative mining method to maximize the utilization of high-quality negative examples. However, real-world datasets often suffer from incomplete positive sample labeling due to cost constraints; for most datasets, only around one positive sample is labeled per query. This necessitates careful consideration of the influence of false negatives during negative sample mining. Many researchers suggest that sampling from the middle of the embedding retrieval results or filtering based on similarity scores can effectively enhance model performance [89, 90]. (3) Pre-training paradigm. Contrastive learning is useful on large-scale corpora but takes big performance hits on low data situations [74]. This observation shows that the pre-training tasks of PLMs, i.e., Masked Language Modeling (MLM) and Next Sentence Prediction (NSP), etc., do not allow the PLMs to adapt to semantic aggregation quickly and output high-quality text embeddings. Therefore, some work has been devoted to designing better pre-training tasks to improve the performance of contrastive learning in low-data scenarios. Early explorations [91, 92] used only additional incremental pre-training tasks, such as the Inverse Cloze Task (ICT), to improve contrast learning performance. Subsequent improvements involve the modification of model architectures. For example, Condenser [93] restores the masked token using the [CLS]’s hidden states from the last layer and other tokens’ hidden states from a shallow layer. Therefore, the newly generated information in the deep layers has to aggregate in the [CLS] to reconstruct the masked information. SEED-Encoder [94] and RetroMAE [95] feed the final hidden state of [CLS] into an auxiliary decoder to restore the masked information, encouraging the aggregation of information at [CLS].
@@ 2.3.1 Semantic Textual Similarity
Semantic Textual Similarity (STS) task [113] is the task to judge whether the similarity calculated by embeddings matches the ratings from humans. Strictly speaking, STS is not a downstream task with a well-defined application scenario but rather an intrinsic evaluation task purposely presented [114]. The motivation is that the distance between the two texts’ embedding should reflect the degree of semantic similarity between these two texts.
@@
The dataset for evaluating STS can be expressed as $D_{\rm eval}^{\rm STS}=\{(x_{i}^{a},x_{i}^{b},c_{i})\}_{i=1}^{n}$, where $(x_{i}^{a},x_{i}^{b})$ is a text pair and $c_{i}$ is the average similarity between the two texts. The popular evaluation datasets usually contain seven STS datasets for English setting, which contains STS12-16 [113, 115, 116, 117, 118], STS-B [114] and SICK-R [119] in SentEval Benchmark [120] and STS-17 [114] and STS-22 [121] for multi-lingual setting.
@@
For evaluation, all texts $\{x_{i}^{a}\}_{i=1}^{n}\cup\{x_{i}^{b}\}_{i=1}^{n}$ are encoded by the learned embedder. A similarity function, e.g. dot product or cosine similarity, is introduced to measure the semantic similarity between the text pairs in the embedding space. Common evaluation metrics include the Pearson correlation coefficient and Spearman correlation coefficient.
@@
Since different methods lead to large differences in the value range of similarity, and we are concerned with relative rankings rather than absolute values in practice [122]. Therefore, the Spearman correlation coefficient is used by default.
@@ 2.3.2 Information Retrieval
Information retrieval (IR)  [^7] Information retrieval in this survey refers to ad-hoc text retrieval by default. aims to retrieve the most related texts relevant to the query from a large-size candidate set. Modern IR systems involve multiple stages of recalling or re-ranking from different-scale candidate sets. Embedding-based retrieval methods are referred to as dense retrieval, which is often used as a method of recalling or secondary-filtering, placed after sparse retrieval (e.g., BM25 [123]) and before reranking (e.g., the cross-encoder scheme recommended by BERT  [68]).
@@
For efficiency of evaluation, all query $\{q_{i}\}_{i=1}^{n}$ and all candidate texts $\{d_{i}\}_{i=1}^{n}\cup\{d_{j}\}_{j=1}^{N}$ can be encoded as embeddings with the learned embedder in advance. Then, a similarity function is introduced to measure the relevance of each query-candidate embedding pair. For each query, all candidate texts will sorted by the predicted relevance in descending and form an ordered candidate list. The common evaluation metrics include Recall Rate, Accuracy, Mean Reciprocal Rank (MRR), Mean Average Precision (MAP) and Normalized Discounted Cumulative Gain (NDCG).
@@
    where $\text{R}_{q_{i}}^{(1)}$ is the rank of the first relevant text in the ordered candidate list.
@@
  • NDCG considers the rank of the relevant text and suggests situating the more pertinent text at the more superior position. DCG for each $q_{i}$ should be calculated first, which can be written as
@@
    where $r_{i}$ is the graded relevance score for the $i$-th retrieved text in the candidate text list. Then $\text{NDCG}@k$ can be calculated by aggregating the normalized DCG values at a specific rank position:
@@ 2.3.3 Universal Embedding
The MTEB benchmark [126] contains 56 datasets in 7 different types of downstream tasks, including classification, clustering, pair classification, reranking, retrieval, and Semantic Textual Similarity (STS), to assess the degree of generalization of text embedding. Until October 2024, MTEB has the community variants in Chinese [127], French [128], Polish [129], Scandinavian [130], Russian [131], Arabic [132] and Persian [133], and an official multilingual version, i.e. MMTEB  [^8] https://github.com/embeddings-benchmark/mteb/blob/main/docs/mmteb, during developing.
@@ 2.4.2 Embedding Inversion
The main metrics for evaluating embedding inversion include BLEU, ROUGE, Exact-Match, cosine similarity, and Token F1. Please refer to Equation 12 and 13 for Exact-Match and ROUGE-L, separately.
  • BLEU measures n-gram similarity between reconstructed text and original input text. We can compute it in the following:
@@ 4 LLMs as Text Embedder
LLMs commonly adopt encoder-decoder architecture, e.g., T5 [98], or decoder-only architecture, e.g., GPT-3 [166]. After going through pre-training, instruction tuning, and alignment with human preferences [167], LLMs can perform a wide range of tasks. However, these steps do not support obtaining high-quality embeddings from LLMs. To address this issue, exploring methods to obtain text embeddings from LLMs directly is a promising direction. In Table 4, we list 40 LLM-based embedders and their detailed information. Specifically, we focus on their backbone selection, architectural improvement, training methods, and evaluation protocols.
@@ 4.1 Backbone Selection
The LLM-based embedders appearing in Table 4 use numerous open-source LLMs as backbones. Among the encoder-decoder backbones, T5 (GTR) has a dominant position, which is not surprising, as T5 and its variants have been leading innovations in encoder-decoder PLMs. Among decoder-only backbones, the most popular is Mistral (20 / 40), followed by LLaMA (17 / 40) and Qwen (6 / 40). The popularity of the Mistral is strongly connected to its excellent performance: a recent study [168] find that Mistral’s embedding performance will not drop significantly compared to the other LLMs when converting causal attention to bi-directional attention. This characteristic may derive from its pre-training or instruction fine-tuning phase, but no information has been revealed.
@@
The vast majority of LLM-based embedders have a backbone size set to 7B, which, directly speaking, may result from a trade-off between efficiency and performance. However, the only several attempts to use LLM with larger LLMs lead to even worse performance [169, 170]. Therefore, the scaling law of LLM-based embedders need to be explored further.
@@ 4.2.1 Pooling Strategy
PromptEOL [170] OPT, LLaMA Last Causal $\times$ $\times$ Prompt TF / SCL ZS / ICL STS / Clf.
@@
Token Prepending [177] LLaMA, Qwen, Gemma Last Causal $\times$ $\times$ Prompt TF ZS / FT STS / Clf.
BeLLM [178] LLaMA Last Bi $\times$ LoRA Prompt SCL ZS / FT STS / Clf.
@@
PromptReps [184] Mistral, Phi-3, LLaMA Last Causal $\surd$ $\times$ Instruction+Prompt TF / SCL ZS IR
LMORT [185] GPT2, GPT-J Multi-Layer Causal $\surd$ $\times$ - SCL FT IR
@@
RARe [187] LLaMA, Mistral Mean / Last Bi / Causal $\times$ LoRA Instruction+Example SCL ICL IR
@@
O1 Embedder [189] Mistral, LLaMA, Qwen Special Last Causal $\times$ LoRA Instruction SCL+NTP ZS IR
Echo [190] Mistral Partial Mean Causal $\times$ LoRA Instruction+Prompt TF / SCL ZS Uni.
@@
SFR-Mistral [194] Mistral Last Causal $\times$ LoRA Instruction SCL ZS Uni.
LLM2Vec [168] LLaMA, Mistral Mean Bi $\times$ LoRA Instruction MNTP→UCL/SCL ZS Uni.
@@
BGE-ICL [202] Mistral, Gemma Special Last Bi / Causal $\times$ LoRA Instruction+Example SCL ZS / ICL Uni.
@@
LENS [204] Mistral Post Bi $\times$ LoRA Instruction SCL ZS Uni.
@@
  • Special-Token Last Pooling. Some works [195, 196, 180, 184, 183, 182, 181] introduce a special token, such as <EOS>, to obtain embedding in the last position, while the LLM is fine-tuned incrementally, learning to converge semantics at the position of the special token. These methods are marked as “special last” in the “pooling” column of Table 4.
  • Special-Sequence Last Pooling. ChatRetriever [183] and DEBATER [188] add a special token sequence, such as $[\texttt{EMB}_{1}],\cdots,[\texttt{EMB}_{t}]$, at the end of the input text. The authors argue that these $t$ consecutive special tokens serve as a chain of thought that extends and guides the learning space for more effective embeddings. These methods are marked as “special seq last” in the “pooling” column of Table 4.
@@
Some works add complex modules with attention mechanisms before the conventional pooling strategy. NV-Embed [199] introduces an additional attention-based network before mean pooling to post-interact with the hidden states $\mathcal{H}$. The attention-based network consists of a cross-attention block and an MLP, where the key and value matrix used in the cross-attention are learnable and low-rank, and $\mathcal{H}$ is considered as the query. Mistral-SPLADE [186] uses the same prompt to Echo [190] to obtain text embeddings and follow the SPLADE recipe [205] to fine-tune the Mistral-based embedded. LENS [204] further clusters the LLM’s tokens according to their embedding to narrow the sparse representation latitude and reduce information redundancy, achieving better sparsification results.